This on states, solve by completing the square:
m^2 - 1/2m = 7/2

Answer by ewatrrr(24785)

 
Hi,
solve by completing the square:
m^2 - 1/2m = 7/2
(m-1/4)^2 - 1/16 = 56/16
(m-1/4)^2 = 57/16
m - 1/4 = ± sqrt(57)/4
m = 1/4 ± sqrt(57)/4